# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ 5.1. Parahoric Group Scheme
A parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$ is equivalent to a compatible tuple (,~$(\mathcal{E},\widetilde{\mathcal{E}}_{i},\Theta_{i})$.
Now we move to principal bundles on root stacks. Let $\alpha$ be a weight and we suppose that its denominator is $d$. This weight $\alpha$ uniquely determines a representation $\rho:\mu_{d}\rightarrow T$ [BS15, Lemma 2.2.8]. Given a collection of weights $\boldsymbol{\alpha}$ with $\boldsymbol{d}$ the corresponding collection of denominators, they determine a collection of representations $\boldsymbol{\rho}$. Since all weights in $\boldsymbol{\alpha}$ are tame, we obtain a (tame) root stack $\mathscr{X}:=\mathscr{X}_{(X,D,\boldsymbol{d})}$ given by Theorem 4.3.
Proposition 5.5.
@@
We prove the correspondence by discussing transition functions. We first recall some notations. Let $R_{i}$ be the discrete valuation rings given by the generic point $\eta_{i}$ of $D_{i}$. We have
By Lemma 4.6, a $G$-bundle $\mathscr{E}$ of type $\boldsymbol{\rho}$ on $\mathscr{X}$ is equivalent to a compatible tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$, where
$\Theta_{i}:\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
is an isomorphism. Recall that
$\Theta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Theta_{i}(w_{i}).$
We can choose an element $\Delta_{i}\in T(\widehat{\widetilde{R}}_{i})$ such that
$\Delta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Delta_{i}(w_{i}),$
which can be regarded as $\Delta_{i}=w_{i}^{d_{i}\alpha_{i}}$ (see [BS15, (2.2.9.4)]), and define new isomorphisms
$\Theta^{\prime}_{i}:=\Delta^{-1}_{i}\Theta_{i},\quad 1\leq i\leq s,$
and it is easy to check that $\Theta^{\prime}_{i}$ is $\mu_{d_{i}}$-invariant, i.e.
$\Theta^{\prime}_{i}(\gamma_{i}\cdot w_{i})=\Theta^{\prime}_{i}(w_{i}).$
Thus, $\Theta^{\prime}_{i}$ descents to a well-defined element in $G(\widehat{K}_{i})$ under the substitution $z_{i}=w_{i}^{d_{i}}$. Thus, there is a one-to-one correspondence between elements in $G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}$ and elements in $G(\widehat{K}_{i})$.
With a similar approach as [BS15, Theorem 2.3.1], we have
More precisely, given an isomorphism $\Psi:\mathscr{E}_{i}\rightarrow\mathscr{E}_{i}$, we regard it as an element in $G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}$, i.e.
$\rho(\gamma_{i})\Psi_{i}(w_{i})\rho(\gamma_{i})^{-1}=\Psi_{i}(\gamma_{i}\cdot w_{i}).$
Define $\Psi^{\prime}_{i}=\Delta_{i}^{-1}\Psi_{i}\Delta_{i}$, which is $\mu_{d_{i}}$-invariant, i.e.
$\Psi^{\prime}_{i}(\gamma_{i}\cdot w_{i})=\Psi^{\prime}_{i}(w_{i}).$
Thus, $\Psi^{\prime}_{i}(w_{i})$ descents to a well-defined element in $G(\widehat{K}_{i})$ and it is easy to check that it is in $G_{\alpha_{i}}(\widehat{R}_{i})$. Indeed, conjugating by $\Delta_{i}$ and substituting $z_{i}=w_{i}^{d_{i}}$ induces the desired isomorphism.
It is easy to check that the compatbility condition for $G$-bundles on root stacks is equivalent to that for parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors. This finishes the proof of this proposition. ∎
Now we suppose that $X$ is projective. Let $P$ be a parabolic subgroup of $G$. We can define a subgroup scheme $\mathcal{P}_{\boldsymbol{\alpha}}\subseteq\mathcal{G}_{\boldsymbol{\alpha}}$ with respect to a given collection of weights $\boldsymbol{\alpha}$ (see [KSZ24, §4.1]). Given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$, a reduction of structure group is a section $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$, and then we obtain a parahoric $\mathcal{P}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}_{\varsigma}$.
${\mathcal{E}_{\varsigma}}$${\mathcal{E}}$${X}$${\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}}$$\scriptstyle{\varsigma}$
Let $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ be a morphism of group schemes over $X$, where we use the subscript $X$ to emphasize that $\mathbb{G}_{m}$ is regarded as a group scheme over $X$. This morphism $\kappa$ called a character. Then, we obtain a line bundle $L(\varsigma,\kappa):=\kappa_{*}\mathcal{E}_{\varsigma}$. Moreover, since a character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ can be regarded as a character $\chi:P\rightarrow\mathbb{G}_{m}$ [KSZ24, Lemma 4.1], the pairing $\langle\boldsymbol{\alpha},\kappa\rangle$ is defined as
$\langle\boldsymbol{\alpha},\kappa\rangle:=\langle\boldsymbol{\alpha},\chi\rangle=\sum_{i=1}^{s}\langle\alpha_{i},\chi\rangle.$
Definition 5.6 (Definition 4.2 in [KSZ24]).
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. Given a reduction of structure group $\varsigma$ and a character $\kappa$, we define the parahoric degree as
$par\deg\mathcal{E}(\varsigma,\kappa)=\deg L(\varsigma,\kappa)+\langle\boldsymbol{\alpha},\kappa\rangle.$
@@
We remind the reader that we fix an ample line bundle $H$ on $X$ and the degree of $\deg L(\varsigma,\kappa)$ is given with respect to $H$. Given the definition of parahoric degree, we define the $R$-stability condition for parahoric torsors.
@@
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any reduction of structure group $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$,
@@
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$, and denote by $\mathscr{E}$ the corresponding $G$-bundle on $\mathscr{X}$ by Proposition 5.5. The results in [KSZ24] implies the following one-to-one correspondences:
  1. (1)
    characters: $\chi:P\rightarrow\mathbb{G}_{m}$ and $\kappa:P_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ [KSZ24, Lemma 4.1],
  2. (2)
    reductions: $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$ and reductions of $\sigma:\mathscr{X}\rightarrow\mathscr{E}/P$ [KSZ24, Lemma 4.10].
Lemma 5.8.
Under the correspondence of characters and reductions of structure group above, there exists a locally free sheaf $\mathscr{L}$ on $\mathscr{X}$ such that
$\deg_{\mathscr{L}}L(\sigma,\chi)=(d_{1}\dots d_{s})par\deg\mathcal{E}(\varsigma,\kappa).$
@@
[…]
$\mathscr{L}:=\sum_{i=1}^{d_{1}}\mathcal{O}_{\mathscr{X}}(i\widetilde{D}_{1}).$
Borne [Bor07, Theorem 4.3] proved the formula
$\deg_{\mathscr{L}}L(\sigma,\chi)=d_{1}(\deg L(\varsigma,\kappa)+\langle\boldsymbol{\alpha},\kappa\rangle).$
Thus,
$\deg_{\mathscr{L}}L(\sigma,\chi)=d_{1}par\deg\mathcal{E}(\varsigma,\kappa).$
This finishes the proof of this lemma. ∎
Proposition 5.9.
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Denote by $\mathscr{E}$ the corresponding $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$. Then $\mathcal{E}$ is $R$-(semi)stable if and only if $\mathscr{E}$ is $R$-(semi)stable. Thus, the category of $R$-(semi)stable parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ is equivalent to the category of $R$-(semi)stable $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$.
Proof.
The proof of this theorem is the same as that of [KSZ24, Theorem 4.12]. Since the characters $\chi:P\rightarrow\mathbb{G}_{m}$ and reductions of structure group $\sigma:\mathscr{X}\rightarrow\mathscr{E}/P$ are in one-to-one correspondence with characters $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ and reductions $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$ respectively, the theorem follows directly from the equality of the corresponding degree by Lemma 5.8. ∎
@@ 5.3. Logahoric Higgs Torsors and Logahoric Connections
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\vartheta$-compatible reduction of structure group $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$,
@@
Logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors $(\mathcal{E},\vartheta)$ on $X$ are in one-to-one correspondence with logarithmic $G$-Higgs bundles $(\mathscr{E},\theta)$ of type $\boldsymbol{\rho}$ on $\mathscr{X}$. Moreover, this correspondence preserves the stability conditions. Thus, the category of $R$-(semi)stable logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors on $X$ is equivalent to the category of $R$-(semi)stable logarithmic $G$-Higgs bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$.
@@
This proposition is proved in the same way as [KSZ24, Theorem 4.13]. By Proposition 5.5, parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors are in one-to-one correspondence with $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$, and we have to prove such a correspondence also holds for Higgs fields. It is enough to work on local charts $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ and ${\rm Spec}\,R_{i}$, and we follow the same notation as in §4.2. The restriction of the Higgs field $\vartheta_{i}$ on $\mathscr{E}_{i}$ is $\mu_{d_{i}}$-equivariant, i.e.
@@
Define $\theta_{i}:=\Delta_{i}^{-1}\vartheta_{i}\Delta_{i}$, and the new element $\theta_{i}$ is $\mu_{d_{i}}$-invariant, i.e.
$\theta_{i}(\gamma_{i}w_{i})=\theta_{i}(w_{i}).$
Thus, $\theta_{i}(z_{i})$ is well-defined by taking $z_{i}=w_{i}^{d_{i}}$. With the same argument as in Proposition 5.5, the new element $\theta_{i}(z_{i})$ gives a section ${\rm Spec}\,R_{i}\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}D)|_{{\rm Spec}\,R_{i}}$. Then, Higgs fields $\vartheta$ of $\mathcal{E}$ are in one-to-one correspondence with Higgs fields $\theta$ of $\mathscr{E}$. ∎
We define nilpotent logahoric Higgs torsors in the same way as that for $G$-Higgs bundles in §2.2, and we introduce the following categories:
@@
  • ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$: the category of nilpotent logarithmic $G$-Higgs bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$ of exponent $\leq p-1$;
@@
Now we consider logarithmic connections. Chen–Zhu gave the definition of (integrable) connections on $\mathcal{G}$-torsors, where $\mathcal{G}$ is an arbitrary smooth affine group scheme over $X$ [CZ15, Appendix]. We follow their setup and define logarithmic integrable connections on parahoric torsors. Let $\mathcal{V}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. A (integrable) logarithmic connection $\nabla^{\prime}$ on $\mathcal{V}$ is a (integrable) connection $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes_{\mathcal{O}_{X}}\Omega_{X}({\rm log}\,D)$ such that the following diagram commutes
@@
The discussion is similar to that of Higgs bundles (Proposition 5.11). We still work on local charts $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ and ${\rm Spec}\,R_{i}$. The only thing we have to show is how the residues of connections change. Let $(\mathscr{V},\nabla)$ be a logarithmic flat $G$-bundle on $\mathscr{X}$. Recall that the transition function $\Theta_{i}$ changes to $\Theta^{\prime}_{i}=\Delta_{i}^{-1}\Theta_{i}$ (Proposition 5.5), which results in the change of Higgs fields $\theta_{i}=\Delta_{i}^{-1}\vartheta_{i}\Delta_{i}$ (Proposition 5.11). Therefore, we define a new connection $\nabla^{\prime}_{i}(w_{i}):=\Delta_{i}(w_{i})\circ\nabla_{i}(w_{i})$, where the action is the gauge action. Regarding the element $\Delta_{i}(w_{i})$ as $w_{i}^{d_{i}\alpha_{i}}$, we have
@@
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\nabla^{\prime}$-compatible reduction of structure group $\varsigma:X\rightarrow\mathcal{V}/\mathcal{P}_{\boldsymbol{\alpha}}$,
@@
Denote by $\boldsymbol{\rho}$ the collection of corresponding representations of $\boldsymbol{\alpha}$. Let $\boldsymbol{\rho}^{p}$ be the collection of representations given in Remark 4.13 and denote by $\boldsymbol{\alpha}^{\prime}$ the corresponding collection of weights. In conclusion, we obtain the logahoric version of nonabelian Hodge correspondence:
@@
The only thing we have to prove is that if $(X,D)$ is $W_{2}(k)$-liftable, the corresponding pairs $(\mathscr{X},\widetilde{D})$ is also $W_{2}(k)$-liftable, where $\mathscr{X}$ is the root stack determined by the data $(X,D,\boldsymbol{d})$. By the well-known covering lemma of Kawamata [Kaw81, Theorem 17], there exists a covering $\pi^{\prime}:Y\rightarrow X$ such that $Y$ is smooth and $D^{\prime}:=(\pi^{\prime*}D)_{red}$ is a reduced normal crossing divisor on $Y$. Moreover, we can assume that the induced morphism $Y\rightarrow\mathscr{X}$ is surjective and étale [Sim11, Lemma 4.2]. Note that although Kawamata and Simpson stated the results only in characteristic zero, their arguments hold in our case due to the assumption that all integers in $\boldsymbol{d}$ are coprime to $p$. Then it is enough to show that the pair $(Y,D^{\prime})$ is $W_{2}(k)$-liftable. This property has been proved in [XW13, Theorem 1.1] and we also refer the reader to [KS20, Notation 2.16] for the case of curves. ∎
@@
Acknowledgement. The first and third named authors are partially supported by the National Key R and D Program of China 2020YFA0713100, CAS Project for Young Scientists in Basic Research Grant No. YSBR-032. The second named author is supported by Guangdong Basic and Applied Basic Research Foundation 2024A04J0334 and 2024A1515011583.

### v2

@@ 5.1. Parahoric Group Scheme
This line bundle $L(\varsigma_{\eta},\kappa)$ can be extended canonically to a parabolic line bundle on a big open subset $X^{\rm big}$ and denote it by $L(\varsigma,\kappa)$.
@@
By Proposition 5.5, denote by $E$ the $(\Gamma,G)$-bundle on $Y$ corresponding to $\mathcal{E}$. With the same argument as [BS15, Proposition 6.3.1], we have the following correspondences:
  • maximal parabolic subgroups $\mathcal{P}\subseteq(\mathcal{G}_{\boldsymbol{\alpha}})_{\eta}$ and maximal parabolic subgroups $P$ of $G$;
  • characters $\mathcal{P}\rightarrow\mathbb{G}_{m,\eta}$ and characters $P\rightarrow\mathbb{G}_{m}$;
  • generic reductions $\varsigma_{\eta}:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$ and $\Gamma$-equivariant generic reductions $\sigma_{\widetilde{\eta}}:Y_{\widetilde{\eta}}\rightarrow E_{\widetilde{\eta}}/P_{\widetilde{\eta}}$.
Given the above correspondences, the corresponding $\Gamma$-equivariant generic reduction of structure group $\sigma_{\widetilde{\eta}}:Y_{\widetilde{\eta}}\rightarrow E_{\widetilde{\eta}}/P_{\widetilde{\eta}}$ is well-defined on an open subset $Y^{\rm big}\subseteq Y$ by [GW20, Theorem 12.60], where ${\rm codim}(Y\backslash Y^{\rm big})\geq 2$, and denote the corresponding reduction of structure group on $Y^{\rm big}$ by $\sigma$. Then we obtain a $\Gamma$-equivariant line bundle $L(\sigma,\chi)$ on $Y^{\rm big}$. By [Bor07], $L(\sigma,\chi)$ corresponds to a parabolic line bundle on a big open subset $X^{\rm big}:=Y^{\rm big}/\Gamma$, which is the desired parabolic bundle. ∎
@@
As a parabolic line bundle, denote by $par\deg L(\varsigma,\kappa)$ the parabolic degree of $L(\varsigma,\kappa)$. Following the above construction, we define the $R$-stability condition for parahoric torsors as follows:
@@
  • for any maximal parabolic subgroup $\mathcal{P}\subseteq(\mathcal{G}_{\boldsymbol{\alpha}})_{\eta}$;
  • for any antidominant character $\kappa:\mathcal{P}\rightarrow\mathbb{G}_{m,\eta}$ trivial on the center;
  • for any generic reduction of structure group $\varsigma_{\eta}:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$,
@@
Proposition 5.10.
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Denote by $\mathscr{E}$ the corresponding $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$. Then $\mathcal{E}$ is $R$-(semi)stable if and only if $\mathscr{E}$ is $R$-(semi)stable.
@@
Under the isomorphism $\mathscr{X}\cong[Y/\Gamma]$, the proof of Lemma 5.8 implies the following one-to-one correspondences
  • characters: $\chi:P\rightarrow\mathbb{G}_{m}$ and $\kappa:P_{\eta}\rightarrow\mathbb{G}_{m,\eta}$,
  • reductions: $\varsigma_{\eta}:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$ and reductions of $\sigma:\mathscr{X}^{\rm big}\rightarrow(\mathscr{E}|_{\mathscr{X}^{\rm big}})/P$ for some big open substack $\mathscr{X}^{\rm big}$.
Therefore, we only have to show the equality
$\deg L(\sigma,\chi)=par\deg L(\varsigma,\kappa),$
which is given in [Bor07, Theorem 4.3]. ∎
@@ 5.3. Logahoric Higgs Torsors and Logahoric Connections
  • for any maximal parabolic subgroup $\mathcal{P}\subseteq(\mathcal{G}_{\boldsymbol{\alpha}})_{\eta}$;
  • for any antidominant character $\kappa:\mathcal{P}\rightarrow\mathbb{G}_{m,\eta}$ trivial on the center;
  • for any $\vartheta$-compatible reduction of structure group $\varsigma:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$,
@@
Logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors $(\mathcal{E},\vartheta)$ on $X$ are in one-to-one correspondence with logarithmic $G$-Higgs bundles $(\mathscr{E},\theta)$ of type $\boldsymbol{\rho}$ on $\mathscr{X}$. Moreover, this correspondence preserves the stability conditions.
@@
By Proposition 5.5, parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors are in one-to-one correspondence with $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$, and we have to prove such a correspondence also holds for Higgs fields. By Hartog’s lemma, we can work on $X^{\rm big}$ and thus, it is enough to work on the associated tuple $(\mathcal{E}_{0},\mathcal{E}_{i},\Theta^{\prime}_{i})$ (resp. $(\mathscr{E}_{0},\mathscr{E}_{i},\Theta_{i})$) of parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors $\mathcal{E}$ on $X$ (resp. $G$-bundles on $\mathscr{X}$) by Remark 5.6. We follow the same notation as in §4.2 and work on $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ (resp. ${\rm Spec}\,R_{i}$) with local coordinate $w_{i}$ (resp. $z_{i}$).
We start with a Higgs field $\theta$ on $\mathscr{E}$. The restriction of the Higgs field $\theta_{i}$ on $\mathscr{E}_{i}$ is $\mu_{d_{i}}$-equivariant, i.e.
@@
Define $\vartheta_{i}:=\Delta_{i}^{-1}\vartheta_{i}\Delta_{i}$, and the new element $\vartheta_{i}$ is $\mu_{d_{i}}$-invariant, i.e.
$\vartheta_{i}(\gamma_{i}w_{i})=\vartheta_{i}(w_{i}).$
Thus, $\vartheta_{i}(z_{i})$ is well-defined by taking $z_{i}=w_{i}^{d_{i}}$. With the same argument as in Proposition 5.5, the new element $\vartheta_{i}$ gives a section ${\rm Spec}\,R_{i}\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}\,D)|_{{\rm Spec}\,R_{i}}$. Gluing $\vartheta_{i}$ via transition functions $\Theta^{\prime}_{i}$, we obtain a Higgs field $X^{\rm big}\rightarrow\mathcal{E}^{\rm big}(\mathfrak{g})\otimes\Omega_{X^{\rm big}}({\rm log}\,D)$ and thus a Higgs field $\vartheta:X\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}\,D)$ by Hartog’s lemma. This construction also holds in the other direction by Remark 5.7 and thus a one-to-one correspondence.
Based on the correspondence for Higgs fields given above and the correspondence for reductions of structure group given in Proposition 5.10, there is a one-to-one correspondence between $\vartheta$-compatible generic reductions $\varsigma:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$ and $\theta$-compatible reductions $\sigma:\mathscr{X}^{\rm big}\rightarrow(\mathscr{E}|_{\mathscr{X}^{\rm big}})/P$ on some big open substack $\mathscr{X}^{\rm big}$. Therefore, the correspondence preserves the stability conditions. ∎
Note that $\mathcal{E}(\mathfrak{g})$ is a locally free sheaf, of which fibers are isomorphic to $\mathfrak{g}$, and a logarhoric Higgs field is a section $\varsigma:X\rightarrow\mathcal{E}(\mathfrak{g})\otimes_{\mathcal{O}_{X}}\Omega_{X}({\rm log}\,D)$. Therefore, nilpotent logahoric Higgs torsors is defined in the same way as that for $G$-Higgs bundles in §2.2. Then we introduce the following categories:
@@
  • ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$: the category of nilpotent logarithmic $G$-Higgs bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$ of exponent $\leq p-1$.
@@
Now we consider logarithmic connections. Chen–Zhu gave the definition of (integrable) connections on $\mathcal{G}$-torsors, where $\mathcal{G}$ is an arbitrary smooth affine group scheme over $X$ [CZ15, Appendix]. We follow their setup and define logarithmic integrable connections on parahoric torsors. Let $\mathcal{V}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. A(n) (integrable) logarithmic connection $\nabla^{\prime}$ on $\mathcal{V}$ is a(n) (integrable) connection $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes_{\mathcal{O}_{X}}\Omega_{X}({\rm log}\,D)$ compatible with the multiplication of $\mathcal{O}_{\mathcal{V}}$ such that the following diagram commutes
@@
The discussion is similar to that of Higgs bundles (Proposition 5.12) and we only work on local charts $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ and ${\rm Spec}\,R_{i}$. The only thing we have to show is how the residues of connections change. Let $(\mathscr{V},\nabla)$ be a logarithmic flat $G$-bundle on $\mathscr{X}$ and denote by $\nabla_{i}$ the restriction of $\nabla$ to $\mathscr{V}_{i}$. Recall that the transition function $\Theta_{i}$ changes to $\Theta^{\prime}_{i}=\Delta_{i}^{-1}\Theta_{i}$ (Proposition 5.5), which results in the change of Higgs fields $\theta_{i}=\Delta_{i}^{-1}\vartheta_{i}\Delta_{i}$ (Proposition 5.12). Therefore, we define a new connection $\nabla^{\prime}_{i}(w_{i}):=\Delta_{i}(w_{i})\circ\nabla_{i}(w_{i})$, where the action is the gauge action. Regarding the element $\Delta_{i}(w_{i})$ as $w_{i}^{d_{i}\alpha_{i}}$, we have
@@
  • for any maximal parabolic subgroup $\mathcal{P}\subseteq(\mathcal{G}_{\boldsymbol{\alpha}})_{\eta}$;
  • for any antidominant character $\kappa:\mathcal{P}\rightarrow\mathbb{G}_{m,\eta}$ trivial on the center;
  • for any $\vartheta$-compatible reduction of structure group $\varsigma:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$,
@@
Denote by $\boldsymbol{\rho}$ the collection of corresponding representations of $\boldsymbol{\alpha}$. Let $\boldsymbol{\rho}^{p}$ be the collection of representations given in Remark 4.13 and denote by $\boldsymbol{\alpha}^{\prime}$ the corresponding collection of weights. In conclusion, we obtain a logahoric version of the nonabelian Hodge correspondence:
@@
The only thing we have to prove is that if $(X,D)$ is $W_{2}(k)$-liftable, the corresponding pairs $(\mathscr{X},\widetilde{D})$ is also $W_{2}(k)$-liftable, where $\mathscr{X}$ is the root stack determined by the data $(X,D,\boldsymbol{d})$. By the Kawamata covering lemma [Kaw81, Theorem 17], there exists a covering $\pi^{\prime}:Y\rightarrow X$ such that $Y$ is smooth and $D^{\prime}:=(\pi^{\prime*}D)_{red}$ is a reduced normal crossing divisor on $Y$. Moreover, we can assume that the induced morphism $Y\rightarrow\mathscr{X}$ is surjective and étale [Sim11, Lemma 4.2]. Note that although Kawamata and Simpson stated the results only in characteristic zero, their arguments hold in our case due to the assumption that all integers in $\boldsymbol{d}$ are coprime to $p$. Then it is enough to show that the pair $(Y,D^{\prime})$ is $W_{2}(k)$-liftable. This property has been proved in [XW13, Theorem 1.1] and we also refer the reader to [KS20, Notation 2.16] for the case of curves. ∎
@@
Acknowledgement. The first and third named authors are partially supported by the National Key R and D Program of China 2020YFA0713100, CAS Project for Young Scientists in Basic Research Grant No. YSBR-032. The second named author is supported by Guangdong Basic and Applied Basic Research Foundation 2024A1515011583 and National Key R and D Program of China 2022YFA1006600.